OpenStudy (anonymous):

simplify the radical expression (8+ √11)(8- √11)

OpenStudy (jdoe0001):

hint: \(\bf (a-b)(a+b) = a^2-b^2\)

jagr2713 (jagr2713):

Since −8√11 and 8√11are like terms, subtract 8√11 from −8√11 to get 0. (64−11√^2) is that right @jdoe0001

OpenStudy (isaiah.feynman):

Way to go @jdoe0001

OpenStudy (anonymous):

thats not an option, but 64 + √11

jagr2713 (jagr2713):

then its that one

OpenStudy (jdoe0001):

hmmm

OpenStudy (isaiah.feynman):

@ashleychase1 look at jdoe0001's terms, they look alike to yours.

OpenStudy (jdoe0001):

\(\bf ({\color{brown}{ a}}-{\color{blue}{ b}})({\color{brown}{ a}}+{\color{blue}{ b}}) = {\color{brown}{ a}}^2-{\color{blue}{ b}}^2 \\ \quad \\ \quad \\ ({\color{brown}{ 8}}+{\color{blue}{ \sqrt{11}}}) ({\color{brown}{ 8}}-{\color{blue}{ \sqrt{11}}})\implies ?\)